OpenStudy (exoexoexo):

the basketball team of martin luther king high school won 12 of its first 15 games. it then won its next 3 games. its winning percentage was then? Can somebody explain to me what this question is asking?

Vocaloid (vocaloid):

percentage of wins = 100*(wins/total games) first off, the team won 12 games + 3 more, so how many games did it win?

OpenStudy (exoexoexo):

@Vocaloid 15 games

Vocaloid (vocaloid):

right, next step... total games = 15 games + 3 more = ?

OpenStudy (exoexoexo):

18 games

Vocaloid (vocaloid):

right, so percentage of wins = 100*(wins/total games) = 100*(15/18) = your answer

OpenStudy (exoexoexo):

about 83%

Vocaloid (vocaloid):

right